OpenStudy (anonymous):

How do you find oscillations per seconds in V = 156sin(120πt)?

OpenStudy (ghazi):

wt= 120 *pi*t therefore w= 120.pi=120*3.14 and omega (w)= 2*3.4*frequency... therefore 120*3.14=2*3.14*frequency

OpenStudy (ghazi):

can you solve this?

OpenStudy (anonymous):

Yes I did thank you so much!

OpenStudy (ghazi):

:)